OpenStudy (osanseviero):

So I have two lines and their equations, I know they intersect in a point A. I need to find the coordinates

OpenStudy (osanseviero):

r1= (10 6 -1) + s(2 -5 -2) r2= (2 1 -3) + t(3 5 2)

OpenStudy (osanseviero):

So r1=r2, but I am having problems with solving it

OpenStudy (anonymous):

I got you solution ;) Just hold on.

OpenStudy (osanseviero):

What I did: 10 + 2s= 2 + 3t 6-5s = 1 +5t -1-2s=-3+2t So I added 1 and 3 and got: 9=-1 + 5t So t=10 And with that I get that s= -9 But it is not working for the first equation

OpenStudy (osanseviero):

Ok...and why isn't my method working?

OpenStudy (osanseviero):

I can put step by step of what I did if you want

OpenStudy (osanseviero):

The problem is that I am not sure I should use reduce form

OpenStudy (anonymous):

It's better use reduce form. I'll see what you did.

OpenStudy (osanseviero):

Thanks

OpenStudy (osanseviero):

I think I did a dumb error...give me a sec

OpenStudy (osanseviero):

Ok...if I add 1 and 3 I get 9=5t-1 So t=2 If I substitute t in eq 2 I get 6-5s=1+10 -5s=5 s=-1

OpenStudy (anonymous):

Ops, i find an error here. \[r_1\] \[x=10+2s\]\[y=6-5s\]\[z=-1-2s\] \[r_2\] \[x=2+3t\]\[y=1+5t\]\[z=-3+2t\] chose 2 equation \[10+2s=2+3t\]\[6-5s=1+5t\] \[t=2~~and~~s=-1\] now replace this values \[r_1\] \[x=10+2*(-1)\]\[y=6-5*(-1)\]\[z=-1-2*(-1)\] \[r_2\] \[x=2+3*2\]\[y=1+5*2\]\[z=-3+2*2\] then we have \[r_1\] \[x=8\]\[y=11\]\[z=1\] \[r_2\] \[x=8\]\[y=11\]\[z=1\]

OpenStudy (osanseviero):

And that gives (8 11 1) which seems correct for both...but it is different

OpenStudy (osanseviero):

oh...same result :D

OpenStudy (anonymous):

Yeah ;)

OpenStudy (osanseviero):

why was your answer different? (the first one)

OpenStudy (anonymous):

I don't know, I guess that I'm made some mistake.

OpenStudy (osanseviero):

ok :) my mistake was really silly...thanks a lot

OpenStudy (anonymous):

You're welcome